OpenStudy (jigggaaaa74):

what is a factor of x^2-x-12

hartnn (hartnn):

can you find 2 numbers with product = -12 and sum = -1 ?

OpenStudy (himansh):

here the 2 no will be -4 and 3 since product of these two no is -12 and sum of these two no is -1 x^2-x- 12=x^2-4x+3x -12 = x(x-4)+ 3(x-4) =(x-4)(x-3)